A Yo-Yo with a diameter of 0.10 m with a mass of 0.2 kg spins at 5000 rpms with a Torque of 40 Nm. What would its Angular Momentum be?


1
Expert's answer
2021-04-28T13:04:22-0400

Find the moment of inertia:


"I=\frac12 mr^2."

The angular momentum is


"L=I\omega=\frac12 mr^2 \omega=\frac12 mr^2 ·2\pi n=mn\pi r^2,\\\space\\
L=0.2·(5000/60)·3.14·0.05^2=0.13\text{ kg}·\text{m}^2/\text{s}."